OpenStudy (kkutie7):

Write out the first four nonzero terms of the Taylor series for f(x) = 5/(1 − x) about 0. Simplify all coefficients. *my work will be written below shortly* I just want to know why I got it wrong.

OpenStudy (kkutie7):

f(x)=5/(1-x) f(0)=5 f'(x)=5/((1-x)^2) f'(0)=5 f''(x)=5/((1-x)^3) f''(0)=10 f'''(x)=5/((1-x)^4) f'''(0)=30 f''''(x)=5/((1-x)^5) f''''(0)=120

OpenStudy (kkutie7):

5+5x+10x^2/2+30x^3/6+120x^4/24 5+5x+5x^2+5x^3+5x^4

OpenStudy (freckles):

\[\frac{d}{dx}5(1-x)^{-2} \\ =5(-2)(1-x)^{-3}(-1) \\ =10(1-x)^{-3} \\ =\frac{10}{(1-x)^{3}}\]

OpenStudy (freckles):

looks like you did the same thing on the other higher order derivatives

OpenStudy (kkutie7):

Oh crap I didn't write that in =(, but I did make up for it in f''(0) and f'''(0)...

OpenStudy (freckles):

that is interesting how did you find f''(0) without the correct f''(x) :p

OpenStudy (kkutie7):

I have it written correctly on the note pad I have currently in m lap haha. It just didn't translate on to here for some reason.

OpenStudy (freckles):

and yeah your evaluations at x=0 look right

OpenStudy (kkutie7):

Should have been: f(x)=5/(1-x) f(0)=5 f'(x)=5/((1-x)^2) f'(0)=5 f''(x)=10/((1-x)^3) f''(0)=10 f'''(x)=30/((1-x)^4) f'''(0)=30 f''''(x)=120/((1-x)^5) f''''(0)=120

OpenStudy (freckles):

The only thing I can see wrong in your answer is you have 5 terms instead of 4

OpenStudy (kkutie7):

well I started at zero because I tried to start at the first term and I got the answer wrong.

OpenStudy (freckles):

5+5x+5x^2+5x^3+5x^4 I was talking instead of doing 5 terms the question said do 4 so it should be 5+5x+5x^2+5x^3

OpenStudy (kkutie7):

oh I didn't think of it like that. do you mind working on a different on with me real quick?

OpenStudy (freckles):

i can try

OpenStudy (kkutie7):

Find the first four nonzero terms of the Taylor series for 1/y about y = −3. Simplify all coefficients.

OpenStudy (freckles):

f''(-3)=-2/27

OpenStudy (freckles):

f'''(-3)=-6/81

OpenStudy (kkutie7):

oops

OpenStudy (freckles):

but yeah you knew that you had your derivatives right

OpenStudy (freckles):

sorta lol

OpenStudy (freckles):

didn't know you said f'''(x)=-6/x^5 you meant f'''(x)=-6/x^4

OpenStudy (kkutie7):

f(y)=1/y, f(-3)=-(1/3) f'(y)=-(1/y^2), f'(-3)=-(1/9) f''(y)=2/(y^-3), f''(-3)=2/27 f'''(y)=-(6/(y^4)), f'''(-3)=-(6/81)

OpenStudy (freckles):

there is another type-o but I'm just going to shutup now :p

OpenStudy (freckles):

f(y)=1/y, f(-3)=-(1/3) f'(y)=-(1/y^2), f'(-3)=-(1/9) f''(y)=2/(y^3), f''(-3)=-2/27 f'''(y)=-(6/(y^4)), f'''(-3)=-(6/81) there all better and then it is the plug in time

OpenStudy (freckles):

\[f(a)+f'(a)(x-a)+f''(a)\frac{(x-a)^2}{2}+f'''(a) \frac{(x-a)^3}{3 \cdot 2 }\]

OpenStudy (freckles):

where a=-3

OpenStudy (kkutie7):

-1/3+((-1/9)(x+3))+(((-2/27)(x+3)^2)/2)+(((-6/81)(x+3)^3)/6)

OpenStudy (freckles):

and a little simplifying can be done

OpenStudy (kkutie7):

-1/3+((-1/9)(x+3))+(((-1/27)(x+3)^2))+(((-1/81)(x+3)^3))

OpenStudy (freckles):

looks pretty

OpenStudy (freckles):

pretty ugly :p

OpenStudy (freckles):

but I mean correct

OpenStudy (kkutie7):

haha I'm going to check and see if I get it right =)

OpenStudy (freckles):

go you!

OpenStudy (kkutie7):

wrong D=

OpenStudy (freckles):

no way!

OpenStudy (freckles):

\[\frac{-1}{3}-\frac{1}{9}(x+3)-\frac{1}{27}(x+3)^2-\frac{1}{81}(x+3)^3\] you typed in this?

OpenStudy (freckles):

i mean if so can i see a picture of it

OpenStudy (freckles):

oh darn

OpenStudy (freckles):

oh we are wrong

OpenStudy (kkutie7):

maybe I should try terms 1-4 excluding 0?

OpenStudy (freckles):

we forgot that we were dealing with x not y

OpenStudy (freckles):

i mean y not x

OpenStudy (freckles):

those dang x's need to be y's I'm sorry

OpenStudy (kkutie7):

OH! i didn't even notice

OpenStudy (freckles):

yeah and I totally forgot too

OpenStudy (kkutie7):

Yeah that was it lol

OpenStudy (freckles):

did you still get full credit?

OpenStudy (kkutie7):

Yes I did thanks

OpenStudy (freckles):

k cool stuff